OpenStudy (twizttiez):

2. Celeste made 6 cone-shaped ice blocks that each have a radius of 2.8 in. and a height of 7 in. She used a cylindrical container with a radius of 9 in. to fill all of the cones. What is the height of the cylindrical container? Use 3.14 to approximate pi and round your answer to two decimal places. Show your work.

OpenStudy (twizttiez):

@campbell_st

OpenStudy (twizttiez):

@mathmath333

OpenStudy (twizttiez):

@confluxepic

OpenStudy (twizttiez):

I just need someone to walk me through the steps is all

OpenStudy (twizttiez):

@sleepyjess

OpenStudy (twizttiez):

\(\pi r^2 \dfrac h3\)

OpenStudy (twizttiez):

\(\pi r^2 h\)

OpenStudy (twizttiez):

Okay one second ^-^

OpenStudy (twizttiez):

Okay so for the cones I got 160.83498667

OpenStudy (twizttiez):

Oops I cubed it

OpenStudy (twizttiez):

One second lemme redo it

OpenStudy (twizttiez):

Okay so for the cones I got 57.44106667

OpenStudy (twizttiez):

Oh well I don't know the height of the cylinder how can I find that?

OpenStudy (sleepyjess):

hopefully this goes through 57.441..... is correct, for the cylinder, we know that the radius is 9, so \(\pi 9^2 h\), that has to equal the cone, now we square 9 and \(\pi\) is 3.14, so \(3.14 *81 h = 57.016667\), now we solve for h

OpenStudy (twizttiez):

Okay that makes sense

OpenStudy (twizttiez):

But how?

OpenStudy (sleepyjess):

Can you solve from there? :)

OpenStudy (sleepyjess):

Well, first what is 3.14*81?

OpenStudy (twizttiez):

Ohhhhhhh okaiii ^-^

OpenStudy (twizttiez):

254.34?

OpenStudy (sleepyjess):

Then once you find that, divide the 57.01..... by that number yep :)

OpenStudy (twizttiez):

57.44106667/254.34=0.22584362?

OpenStudy (twizttiez):

@sleepyjess

OpenStudy (twizttiez):

Well I'm guessing that 0.22584362 is my height so I plugged it in and got 57.44

OpenStudy (twizttiez):

Thanks!